OpenStudy (anonymous):

What two numbers add to -3 and multiply to 2

OpenStudy (raden):

-1and -2

OpenStudy (raden):

-1 * -2 = 2 -1 + (-2) = -3

OpenStudy (anonymous):

algebraically, let \[x + y = -3\] and \[xy = 2\] solve by substitution.